# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Abstract
Achieving high-quality solutions faster than classical solvers on computationally hard problems is a challenge for quantum optimization to deliver utility. Using a superconducting quantum computer, we experimentally investigate the performance of a hybrid quantum-classical algorithm inspired by semidefinite programming approaches for solving the maximum cut problem on 3-regular graphs up to several thousand variables. We leverage the structure of the input problems to address sizes beyond what current quantum machines can naively handle. We attain an average performance of 99% over a random ensemble of thousands of problem instances. We benchmark the quantum solver against similarly high-performing classical heuristics, including the Gurobi optimizer, simulated annealing, and the Burer-Monteiro algorithm. A runtime analysis shows that the quantum solver on large-scale problems is competitive against Gurobi but short of others. We explore multiple leads to close the gap and discuss prospects for a practical quantum speedup.
@@ I Introduction
The relatively small number of quantum bits (qubits) and high error rates in logical quantum operations of contemporary quantum hardware diminish the practical implementation of such algorithms to solve relevant problems. Nevertheless, experimental realizations are catching up, enabling the co-design of more efficient and performant quantum solvers, unleashing a quantum potential for combinatorial optimization. For instance, a scaling advantage was observed for solving maximum independent set problems up to 289 variables using individual atoms trapped in optical tweezers serving as qubits [6]. Quantum annealers have also been found to deliver a scaling advantage for solving spin glass problems with up to 5,000 variables [7]. Although a scaling speedup paves the way for a possible quantum advantage in runtime, achieving a practical runtime speedup against state-of-the-art classical solvers requires additional theoretical and technological developments.
@@
Here, we use a gate-based superconducting quantum computer featuring transmon qubits [8, 9] to investigate quantum optimization algorithms for solving maximum cut problems up to several thousand variables. The maximum cut problem seeks the bipartition of a graph such that the number of edges connecting vertices from the two distinct groups, which we call the cut number, is maximized. Devising an algorithm that can guarantee a cut number at least $94.1\%$ of the optimal one on an arbitrary graph belongs to the computational complexity class NP-hard [10, 11]. To date, the algorithm guaranteeing the highest cut number, at $87.9\%$ of the optimal one, comes from Goemans and Williamson [12]. In practice, heuristic solvers with limited or no guarantees are used, as they usually yield solutions above those theoretical thresholds and often faster. Therefore, the main competitors for a practical utility are commercial solvers such as the Gurobi optimizer [13], state-of-the-art heuristics such as simulated annealing [14], and methods tailored towards specific applications like the Burer-Monteiro approach for the maximum cut problem [15, 16]. Quantum computers have been used to solve maximum cut problems [17, 18, 19, 20], but their results have lagged behind classical solvers in the performance versus absolute runtime space to warrant a direct benchmark. Here, we pave the way for a practical quantum speedup by achieving an experimental cut number at $99\%$ of optimal up to several thousand variables in tens of seconds on a class of problems.
II The Maximum Cut Problem on a Quantum Computer
@@
Figure 1: (a) Embedding of a light-cone-induced $7$-vertex subgraph by the QAOA at $p=1$ for a $3$-regular graph onto a linear chain of $7$ qubits. (b) Experimental performance distribution of the greedy-enhanced quantum relax-and-round (QRR*) solver with $p=1$ over $1,000$ randomly generated problem instances for various number of variables. (c) Average experimental performance as a function of the problem size for the quantum approximate optimization algorithm (QAOA), the quantum relax-and-round (QRR), and the QRR* solvers with $p=1$. (d) Simulated average performance of the QRR* solver as a function of the number of variables for $p=1$ and $p=2$ layers. Error bars indicate one standard deviation.
@@
defined over binary variables $z_{i}=\pm 1$ assigned to each of the $N$ graph vertices and labeling their bipartition assignment. The graph structure is encoded in its adjacency matrix $\mathsf{W}$ where $\mathsf{W}_{ij}=1$ if vertices $i$ and $j$ are connected by an edge and zero otherwise. Denoting $\boldsymbol{z}_{\textbf{opt}}$ the optimal solution, $\alpha=C(\boldsymbol{z})/C(\boldsymbol{z}_{\textbf{opt}})$, known as the approximation ratio, measures the quality of a solution $\boldsymbol{z}$.
A quantum algorithm will typically formulate an objective function, such as Eq. (1), as a quantum many-body Hamiltonian by mapping each of the variables onto a qubit [21]. Quantum physical principles like adiabatic time evolution [2, 4] or variational search [3, 5] are then employed to find its ground state. In the context of gate-based quantum computing, the quantum approximate optimization algorithm (QAOA) [3, 5] based on such a variational search and parametrized by its number of layers $p$, is particularly well-suited [22, 23, 17]. However, the qubit count and error rate in current quantum devices make the QAOA uncompetitive in practice [17, 24, 25, 26].
To overcome those limitations, we embed the QAOA into a relax-and-round approach [27], which we complement with a guided greedy local search. We solve the maximum cut problem on random $3$-regular graphs in which each vertex is connected at random to exactly three others. Thanks to the input graph structure and the relax-and-round strategy utilizing expectation values, we can leverage a light cone technique to solve problem sizes with thousands of variables, exceeding what today’s quantum machines can naively handle. We achieve an average approximation ratio of $99\%$ across thousands of problem instances. For reference, on such graphs, a perfect implementation of the QAOA at $p=1$ guarantees a cut of $69.2\%$ of the optimal one [3, 28].
@@
where $\hat{H}$ is the one-qubit Hadamard gate, $\hat{X}_{i}$ is the Pauli operator on qubit $i$, and $\langle\boldsymbol{z}|\hat{C}|\boldsymbol{z^{\prime}}\rangle=C(\boldsymbol{z})\delta_{\boldsymbol{z}\boldsymbol{z^{\prime}}}$ a diagonal unitary encoding the objective function. The parameters $\gamma_{\ell}$ and $\beta_{\ell}$ are variational and optimal when maximizing the expectation value of the objective function. Instead of considering sampled solutions $\{\boldsymbol{z}\}$ from the quantum computer directly, the quantum relax-and-round (QRR) algorithm [27] works with the expectation value of two-body correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ where $\hat{Z}_{i}$ is the Pauli operator on qubit $i$. Working with expectation values opens the door to a wide range of error mitigation techniques [29, 30] to be applied to optimization problems. From the point of view of qubit $i$, a $p$-layer QAOA circuit only induces quantum coherence with its $p$-nearest neighbors according to the graph topology. This collection of coherent qubits around $i$ is known as its light cone. To evaluate a two-body expectation value between two qubits $i$ and $j$, one only needs to consider the quantum circuit running on the subset of qubits made of the union of their respective light cones [3]. The correlation of interest is non-zero only if the light cones intersect, leading to only $O(N)$ non-trivial correlations when $p\ll\ln N$ for $N$-variable $3$-regular graphs. Moreover, the sparse nature of $3$-regular graphs is such that subcircuits involve at most $1+6(2^{p}-1)$ qubits, independently of $N$. Therefore, one can trade a single QAOA execution on $N$ qubits for $O(N)$ smaller executions and access graph problem sizes much larger than current hardware would normally support [30].
@@ III Quantum Optimization
For the QAOA with $p=1$ on $3$-regular graphs, one needs seven qubits at most (Fig. 1a). We use transmon qubits [8, 9] arranged linearly. We compile the original quantum logical operations of Eq. (2) at the individual qubit level through single-qubit rotation gates Rz (virtual, error-free) and Rx (average error rate of $0.8\%$[30]. We realize entangling operations, mediated by a tunable coupler [31, 32] between nearest-neighboring qubits, in the form of ISWAP gates (average error rate of $1.9\%$ [30]). The original quantum logical operations of Eq. (2) between non-neighboring qubits are implemented via an extensive swap network [33, 30]. We employ readout error mitigation [34, 30]. We use fixed values for the variational parameters $\gamma_{\ell}$ and $\beta_{\ell}$, expected to produce a decent average performance on any $3$-regular graph, albeit sub-optimal on a case-by-case basis [35, 36, 30]. It is a trade-off eliminating the need for variational optimization of the quantum circuit parameters [37]. We have found that collecting $n\gtrsim 5\times 10^{3}$ bit strings from the quantum computer for each observable yields the best performance [30]. In practice, we use $n=10^{4}$.
@@
Once two-body correlations are collected for a given problem instance, the QRR algorithm requires the construction of a correlation matrix with entries $\mathsf{Z}_{ij}=(\delta_{ij}-1)\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$. A solution to the maximum cut problem is then obtained by sign-rounding one of its leading eigenvectors [30]. In the guided greedy local search that we use to augment the relax-and-round approach, instead of sign-rounding the selected eigenvector directly, we interpret the magnitude of its entries as a confidence measure of their sign. We leverage this information to implement a greedy local step that will revisit variables with low-confidence values. On such values, a sign-flip is attempted and retained as part of the solution if it improves the cut number [30]. We refer to this variant of the solver as the greedy-enhanced quantum relax-and-round (QRR*) solver. Its performance distribution across thousands of problem instances from sizes $N=32$ to $N=4,096$, and based on the QAOA with $p=1$, is displayed in Fig. 1b. We achieve an average approximation ratio of $99\%$. For problem sizes up to $N\leq 256$, we find the optimal solution for a finite fraction of the problems. Although the light-cone technique does not enable us to sample solutions from the underlying QAOA directly, we can estimate its expected performance by computing the expectation value of the objective function $\langle\hat{C}\rangle_{p}$. We report the corresponding average approximation ratio over all problem instances in Fig. 1c. We find that the QRR algorithm enhances the average performance, relative to QAOA, by an order of magnitude using the same underlying quantum resources, ($1-\alpha\simeq 30\%\to 3\%$). We observe that augmenting QRR with a local greedy search (QRR*) improves the average performance by a factor of three, ($1-\alpha\simeq 3\%\to 1\%$).
@@
At $p=1$, the quantum algorithm gives the same result as a classical relax-and-round approach performed on the adjacency matrix of the graph problem $\mathsf{W}$, rather than the correlation matrix $\mathsf{Z}$. Indeed, one can show that they approximately share the same eigenvectors (and presumably the same leading eigenvectors), resulting in the same cut number [30]. By increasing the number of QAOA layers $p$, the classical and quantum relax-and-round differentiate, with QRR converging asymptotically with $p$ to the optimal solution [27, 30]. In Fig. 1d, we show the average performance of the QRR* algorithm from classically emulated quantum circuits at $p=1$ and $p=2$, displaying a performance increase with $p$. The gap reduction in performance between the two is a finite-size effect similarly observed in the QAOA [30]. Nonetheless, QRR* provides a very high-performance baseline, matching that of a classical algorithm (even with minimal quantum resources, $p=1$), and also a path towards improving the solution quality by increasing $p$.
@@ IV Quantum Speedup
Figure 2: Optimal runtime for various classical solvers to match the performance of the greedy-enhanced quantum relax-and-round (QRR*) solver with $p=1$, normalized by the runtime of QRR* [30], as a function of the number of variables. The Gurobi data points for the two largest sizes are lower bounds where no solutions matching that of QRR* were found in $600$ seconds for most of the problem instances [30]. The quantum runtime component of QRR* is based on $t_{\texttt{Rx}}=40$ ns, $t_{\texttt{ISWAP}}=122$ ns, $t_{\texttt{m+r}}=6~{}\mu$s, and $n=5\times 10^{4}$ shots [30]. (a) Utilizing a quantum computer with $10^{2}$ qubits. (b) Utilizing a quantum computer with $10^{3}$ qubits. Each data point is averaged over $1,000$ randomly generated problem instances for each number of variables. Error bars indicate one standard deviation.
The relatively high performance of QRR* makes its solution meaningful for the maximum cut problem on $3$-regular graphs. We turn our attention to a benchmark against state-of-the-art classical solvers and focus on three main competitors (as well as others [30]): simulated annealing (SA) [14], Gurobi [13], and the Burer-Monteiro (BM) algorithm [15, 16]. The first two are general-purpose solvers for combinatorial optimization, while BM is tailored for the maximum cut problem and currently the best algorithm for solving $3$-regular graphs [38]. Whereas QRR* is also not restricted to the maximum cut problem, skipping the variational search by using fixed parameters [35, 36, 30] makes it special-purpose. Neither Gurobi, nor BM, have hyperparameters that need to be optimized for the solver to work best. On the other hand, SA requires a temperature schedule. Given that the quantum solver has no hyperparameters, we rely on a constant-time heuristic setting of the temperature schedule [30].
A classical solver typically performs better given more runtime. Runtime is the only parameter given to solvers like Gurobi and BM while SA runtime is directly proportional to the number of sweeps (set by the user) [30]. The control parameter of the quantum algorithm is the number of layers $p$. Because imperfect quantum machines limit the practical size of quantum circuits that can be executed, one is constrained to a low-$p$ regime. All of the classical solvers tested here are randomized, thus returning a different solution for each run. We define $*P(C(\boldsymbol{z}_{\textrm{cl}})\geq C(\boldsymbol{z}_{\textrm{QRR*}}))$ to be the probability for a classical solver with runtime $t$ to return a solution $\boldsymbol{z}_{\textrm{cl}}$ at least matching that of the quantum algorithm $*\boldsymbol{z}_{\textrm{QRR*}}$, so that
@@
is the average time for state-of-the-art classical solvers to match the performance of the QRR* algorithm. $P$ is intrinsically related to $t$ given that a longer runtime is more likely to return a better solution, and conversely, a smaller runtime $t$ is more likely to result in a lower value of $P$. Therefore we seek the optimal time $t^{*}_{\textrm{opt}}$ for the classical solver to find a solution matching the quality of the QRR* solution, which concurrently minimizes $t$ and maximizes $P$. We compute on a per-problem-instance basis $t^{*}_{\textrm{opt}}$ for each of the three classical solvers [30].
@@
We compare $t^{*}$ with the runtime for QRR*, which we estimate as a function of the problem size $N$, the number of QAOA layers $p$, the total number of qubits available $M$ on a device, and the number of bit strings sampled $n$ [30]. There are four components to consider: executing the quantum circuits on the quantum computer, building the correlation matrix, finding its leading eigenvectors, and doing the greedy local search. The last three steps are purely classical. The execution time is based on the use of the light-cone technique on an $M$-qubit quantum computer. Correlation-induced subcircuits by the QAOA with $p$ layers on an $N$-variable $3$-regular graph problem are executed simultaneously, which may require multiple sequential runs if $M$ is too small to encapsulate them all at once [30]. The execution time of the subcircuits is based on their compilation onto a linear chain topology of qubits using a worst-case-scenario swap network [33] with hardware-native gates ($t_{\texttt{Rx}}=40$ ns and $t_{\texttt{ISWAP}}=122$ ns), measurement, and reset times ($t_{\texttt{m+r}}=6~{}\mu$s), all multiplied by $n=5\times 10^{4}$ (shot count) [30].
@@
We consider two scenarios where a quantum computer has $M=10^{2}$ (Fig. 2a) and $10^{3}$ (Fig. 2b) qubits, enabling the parallel execution of more QAOA subcircuits. This range corresponds to what is already and should be available within the next few years, respectively. We show the optimal time $t^{*}_{\textrm{opt}}$ for classical solvers to match the experimental performance of QRR* with $p=1$, normalized by the runtime of QRR* itself. In the case of Gurobi, this quantum speedup factor exceeds one for the largest problem sizes, meaning that QRR* is faster. In fact, for the two largest sizes, we only provide a lower bound corresponding to a ten-minute runtime limit for Gurobi (greyed-out data symbols in Fig. 2), which failed to match the performance of QRR* for most of the $1,000$ problem instances [30]. Regardless, Gurobi provides additional valuable information, such as a bound on how close a solution is to optimality and it can guarantee the optimality of a solution. The SA and BM solvers are faster than QRR* at a fixed number of QAOA layers, with the gap closing as problem sizes increase. For SA, we assumed knowledge of the optimal number of sweeps, which would otherwise need to be searched. Consequently, the quantum speedup factor for SA should be considered a lower bound.
@@
The first path for improving the competitiveness of the quantum solver is to reduce its runtime. The classical steps could be optimized using standard software engineering techniques and/or algorithmic improvements, and one could imagine the control system of the device estimating expectation values on the fly rather than collecting bit strings that require post-processing. The execution time on the quantum computer could also be reduced through a more efficient compilation of the quantum circuit onto the hardware-native topology of the qubits, shorter logical gates, and the development of new gates, reducing the compilation overhead. The second path for improving competitiveness is to increase the performance of QRR*. While additional classical processing is an option, increasing the number of layers $p$ in quantum algorithms remains the cornerstone of more performant quantum algorithms.

### v2

# Benchmarking Quantum Optimization for the Maximum-Cut Problem on a Superconducting Quantum Computer

@@ Abstract
Achieving high-quality solutions faster than classical solvers on computationally hard problems is a challenge for quantum optimization to deliver utility. Using a superconducting quantum computer, we experimentally investigate the performance of a hybrid quantum-classical algorithm inspired by semidefinite programming approaches for solving the maximum-cut problem on 3-regular graphs up to several thousand variables. We leverage the structure of the input problems to address sizes beyond what current quantum machines can naively handle. We attain an average approximation ratio of 99% over a random ensemble of thousands of problem instances. We benchmark the quantum solver against similarly high-performing classical heuristics, including the Gurobi optimizer, simulated annealing, and the Burer-Monteiro algorithm. A run-time analysis shows that the quantum solver on large-scale problems is competitive against Gurobi but short of others on a projected $100$-qubit quantum computer. We explore multiple leads to close the gap and discuss prospects for a practical quantum speedup.
@@ I Introduction
However, executing these algorithms at scale on a large ensemble of problems to thoroughly benchmark them against classical solvers, and open avenues to outperform the latter, has faced significant challenges. First, the relatively small number of quantum bits (qubits) and high error rates in logical quantum operations of contemporary quantum hardware diminish the practical implementation of such algorithms to solve relevant problems. Nevertheless, experimental realizations are catching up, enabling the codesign of more efficient and performant quantum solvers, unleashing a quantum potential for combinatorial optimization. For instance, a scaling advantage has been observed for solving maximum-independent-set problems up to $289$ variables using individual atoms trapped in optical tweezers serving as qubits [6, 7]. Quantum annealers have also been found to deliver a scaling advantage for solving spin-glass problems with up to 5,000 variables [8]. Second, although a scaling speedup paves the way for a possible asymptotic quantum advantage in run-time, achieving a practical run-time speedup against state-of-the-art classical solvers at relevant scales requires additional theoretical and technological developments.
@@
Here, we use a gate-based superconducting quantum computer featuring transmon qubits [9, 10] to extensively benchmark quantum optimization algorithms for solving thousands of maximum-cut problems with each having up to several thousand variables. The maximum-cut problem seeks the bipartition of a graph such that the number of edges connecting vertices from the two distinct groups, which we call the cut number, is maximized. Devising an algorithm that can guarantee a cut number at least $94.1\%$ of the optimal one on an arbitrary graph belongs to the computational complexity class NP-hard [11, 12]. To date, the algorithm guaranteeing the highest cut number, at $87.9\%$ of the optimal one, comes from Goemans and Williamson [13]. In practice, heuristic solvers with limited or no guarantees are used, as they usually yield solutions above those theoretical thresholds and often faster. Therefore, the main competitors for a practical utility are commercial solvers such as the Gurobi optimizer [14], state-of-the-art heuristics such as simulated annealing [15], and methods tailored toward specific applications, such as the Burer-Monteiro (BM) approach for the maximum-cut problem [16, 17]. Quantum computers have been used to solve maximum-cut problems [18, 19, 20, 21], but their results have lagged behind classical solvers in the performance-versus-absolute-run-time space to warrant a direct benchmark. Here, we pave the way for a practical quantum speedup by achieving an experimental cut number at $99\%$ of optimal up to several thousand variables in tens of seconds on a class of problems.
II The Maximum-Cut Problem on a Quantum Computer
@@
Figure 1: (a) The embedding of a light-cone-induced seven-vertex subgraph by the quantum approximate optimization algorithm (QAOA) at $p=1$ for a $3$-regular graph onto a linear chain of seven qubits. (b) Experimental performance distribution of the greedy-enhanced quantum relax-and-round (QRR*) solver with $p=1$ over $1,000$ randomly generated problem instances for various numbers of variables. (c) The average experimental performance as a function of the problem size for the QAOA, the quantum relax-and-round (QRR), and the QRR* solvers with $p=1$. (d) The simulated average performance of the QRR* solver as a function of the number of variables for $p=1$ and $p=2$ layers. The error bars indicate one standard deviation.
@@
defined over binary variables $z_{i}=\pm 1$ assigned to each of the $N$ graph vertices and labeling their bipartition assignment. The graph structure is encoded in its adjacency matrix $\mathsf{W}$, where $\mathsf{W}_{ij}=1$ if vertices $i$ and $j$ are connected by an edge and zero otherwise. Denoting $\boldsymbol{z}_{\textbf{opt}}$ as the optimal solution, $\alpha=C(\boldsymbol{z})/C(\boldsymbol{z}_{\textbf{opt}})$, known as the approximation ratio, measures the quality of a solution $\boldsymbol{z}$.
A quantum algorithm will typically formulate an objective function, such as Eq. (1), as a quantum many-body Hamiltonian by mapping each of the variables onto a qubit [22]. Quantum physical principles such as adiabatic time evolution [2, 4] or variational search [3, 5] are then employed to find its ground state. In the context of gate-based quantum computing, the quantum approximate optimization algorithm (QAOA) [3, 5], based on such a variational search and parametrized by its number of layers $p$, is one of the most promising [23, 24, 18]. However, even on an ideal quantum computer, the QAOA circuit needs a large depth to outperform classical approximation algorithms for several classes of problems [25, 26],and no practical quantum speedup over state-of-the-art heuristics solvers has been suggested with the QAOA. On actual quantum hardware, the qubit count and error rate make the QAOA uncompetitive in practice [18, 27, 28].
To overcome those limitations, we embed the QAOA into a relax-and-round approach [29], which we complement with a guided greedy local search. We solve the maximum-cut problem on random $3$-regular graphs with thousands of nodes. We achieve an average approximation ratio of $99\%$ across thousands of problem instances. For reference, on such graphs, a perfect implementation of the QAOA at $p=1$ guarantees a cut of only $69.2\%$ of the optimal one [3, 30]. Investigating $3$-regular graphs presents several advantages. First, they are mathematically proven hard to solve [30]. Second, they are sparse, where each vertex is connected randomly to three others. Due to their topology and the relax-and-round strategy utilizing expectation values, we can leverage a light-cone technique to solve problem sizes with thousands of variables, exceeding what today’s quantum machines can naively handle. Third, near-optimal variational QAOA parameters are tabulated [31, 25], significantly speeding up their solving in search of a quantum speedup. Finally, $3$-regular graphs are a standard benchmark problem in quantum optimization since the seminal QAOA work [3].
@@
where $\hat{H}$ is the one-qubit Hadamard gate, $\hat{X}_{i}$ is the Pauli $X$-operator on qubit $i$, and $\langle\boldsymbol{z}|\hat{C}|\boldsymbol{z^{\prime}}\rangle=C(\boldsymbol{z})\delta_{\boldsymbol{z}\boldsymbol{z^{\prime}}}$ is a diagonal unitary encoding the objective function. The parameters $\gamma_{\ell}$ and $\beta_{\ell}$ are variational and optimal when maximizing the expectation value of the objective function. Instead of considering sampled solutions $\{\boldsymbol{z}\}$ from the quantum computer directly, the quantum relax-and-round (QRR) algorithm [29] works with the expectation value of two-body correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$, where $\hat{Z}_{i}$ is the Pauli $Z$-operator on qubit $i$. Working with expectation values opens the door to a wide range of error-mitigation techniques [32] to be applied to optimization problems. From the point of view of qubit $i$, a $p$-layer QAOA circuit only induces quantum coherence with its $p$-nearest neighbors according to the graph topology. This collection of coherent qubits around $i$ is known as its light cone. To evaluate a two-body expectation value between two qubits $i$ and $j$, one only needs to consider the quantum circuit running on the subset of qubits made of the union of their respective light cones [3]. The correlation of interest is nonzero only if the light cones intersect, leading to only $O(N)$ nontrivial correlations when $p\ll\ln N$ for $N$-variable $3$-regular graphs. Moreover, the sparse nature of $3$-regular graphs is such that subcircuits involve at most $1+6(2^{p}-1)$ qubits, independently of $N$. Therefore, one can trade a single QAOA execution on $N$ qubits for $O(N)$ smaller executions and access graph-problem sizes much larger than current hardware would normally support (see App. C.2).
@@ III Quantum Optimization
For the QAOA with $p=1$ on $3$-regular graphs, one needs seven qubits at most (Fig. 1a). We use transmon qubits [9, 10] arranged linearly. We compile the original quantum logical operations of Eq. (2) at the individual qubit level through single-qubit rotation gates Rz (virtual, error-free) and Rx (average error rate of $0.8\%$). We realize entangling operations, mediated by a tunable coupler [33, 34] between nearest-neighboring qubits, in the form of ISWAP gates (average error rate of $1.9\%$). The original quantum logical operations of Eq. (2) between non-neighboring qubits are implemented via an extensive swap network [35]. We employ readout-error mitigation [36]. We use fixed values for the variational parameters $\gamma_{\ell}$ and $\beta_{\ell}$, expected to produce a decent average performance on any $3$-regular graph, albeit suboptimal on a case-by-case basis [31, 25]. It is a trade-off eliminating the need for variational optimization of the quantum circuit parameters [37]. We have found that collecting $n\gtrsim 5\times 10^{3}$ bit strings from the quantum computer for each observable yields the best performance. In practice, we use $n=10^{4}$.
@@
Once two-body correlations are collected for a given problem instance, the QRR algorithm requires the construction of a correlation matrix with entries $\mathsf{Z}_{ij}=(\delta_{ij}-1)\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$. A solution to the maximum-cut problem is then obtained by sign rounding one of its leading eigenvectors. In the guided greedy local search that we use to augment the relax-and-round approach, instead of sign rounding the selected eigenvector directly, we interpret the magnitude of its entries as a confidence measure of their sign. We leverage this information to implement a greedy local step that will revisit variables with low-confidence values. On such values, a sign flip is attempted and retained as part of the solution if it improves the cut number, as described in App. B.2.4. We refer to this variant of the solver as the greedy-enhanced quantum relax-and-round (QRR*) solver. Its performance distribution across thousands of problem instances from sizes $N=32$ to $N=4,096$, and based on the QAOA with $p=1$, is displayed in Fig. 1b. We achieve an average approximation ratio of $99\%$. For problem sizes up to $N\leq 256$, we find the optimal solution for a finite fraction of the problems. Although the light-cone technique does not enable us to sample solutions from the underlying QAOA directly, we can estimate its expected performance by computing the expectation value of the objective function $\langle\hat{C}\rangle_{p}$. We report the corresponding average approximation ratio over all problem instances in Fig. 1c. We find that the QRR algorithm enhances the average performance, relative to the QAOA, by an order of magnitude using the same underlying quantum resources, ($1-\alpha\simeq 30\%\to 3\%$). We observe that augmenting the QRR with a local greedy search (the QRR*) improves the average performance by a factor of $3$, ($1-\alpha\simeq 3\%\to 1\%$).
@@
At $p=1$, the quantum algorithm gives the same result as a classical relax-and-round approach performed on the adjacency matrix of the graph problem $\mathsf{W}$, rather than the correlation matrix $\mathsf{Z}$. Indeed, one can show that they approximately share the same eigenvectors (and presumably the same leading eigenvectors), resulting in the same cut number (Appendix E.4). By increasing the number of QAOA layers $p$, the classical and quantum relax-and-round differentiate, with the QRR converging asymptotically with $p$ to the optimal solution. In Fig. 1d, we show the average performance of the QRR* algorithm from classically emulated quantum circuits at $p=1$ and $p=2$, displaying a performance increase with $p$. The apparent approximation-ratio increase at large $N$ is because we likely underestimate $C(\boldsymbol{z}_{\textbf{opt}})$ by not finding the optimal solution $\boldsymbol{z}_{\textbf{opt}}$; therefore the relative performance of the quantum solver with respect to the best classical solution that we have increases. The gap reduction in performance between the two is a finite-size effect similarly observed in the QAOA (App. E.7). Nonetheless, the QRR* provides a very high-performance baseline, matching that of a classical algorithm (even with minimal quantum resources, $p=1$), and also a path toward improving the solution quality by increasing $p$.
@@ IV Quantum Speedup
The relatively high performance of the QRR* makes its solution meaningful for the maximum-cut problem on $3$-regular graphs. We turn our attention to a benchmark against state-of-the-art classical solvers and focus on three main competitors (as well as others in App. E.7): simulated annealing (SA) [15], Gurobi [14], and the BM algorithm [16, 17]. The first two are general-purpose solvers for combinatorial optimization, while BM is tailored for the maximum-cut problem and is currently the best algorithm for solving $3$-regular graphs [38]. Whereas the QRR* is also not restricted to the maximum-cut problem, skipping the variational search by using fixed parameters [31, 25] makes it special purpose. Neither Gurobi nor the BM have hyperparameters that need to be optimized for the solver to work best. On the other hand, SA requires a temperature schedule. Given that the quantum solver has no hyperparameters, we rely on a constant-time heuristic setting of the temperature schedule (App. B.1.2).
A classical solver typically performs better given more run-time. The run-time is the only parameter given to solvers such as Gurobi and the BM, while the SA run-time is directly proportional to the number of sweeps (set by the user). The control parameter of the quantum algorithm is the number of layers $p$. Because imperfect quantum machines limit the practical size of quantum circuits that can be executed, one is constrained to a low-$p$ regime. All of the classical solvers tested here are randomized, thus returning a different solution for each run. We define $*P(C(\boldsymbol{z}_{\textrm{cl}})\geq C(\boldsymbol{z}_{\textrm{QRR*}}))$ to be the probability for a classical solver with run-time $t$ to return a solution $\boldsymbol{z}_{\textrm{cl}}$ at least matching that of the quantum algorithm $*\boldsymbol{z}_{\textrm{QRR*}}$, so that
@@
is the average time for state-of-the-art classical solvers to match the performance of the QRR* algorithm. $P$ is intrinsically related to $t$ given that a longer run-time is more likely to return a better solution and, conversely, a smaller run-time $t$ is more likely to result in a lower value of $P$. Therefore, we seek the optimal time $t^{*}_{\textrm{opt}}$ for the classical solver to find a solution matching the quality of the QRR* solution, which concurrently minimizes $t$ and maximizes $P$. We compute on a per-problem-instance basis $t^{*}_{\textrm{opt}}$ for each of the three classical solvers (App. E.9).
@@
We compare $t^{*}$ with the run-time for the QRR*, which we estimate as a function of the problem size $N$, the number of QAOA layers $p$, the total number of qubits available $M$ on a device, and the number of bit strings sampled $n$ (Appendix E.8.1). There are four components to consider: executing the quantum circuits on the quantum computer, building the correlation matrix, finding its leading eigenvectors, and doing the greedy local search. The last three steps are purely classical. The execution time is based on the use of the light-cone technique on an $M$-qubit quantum computer. Correlation-induced subcircuits by the QAOA with $p$ layers on an $N$-variable $3$-regular graph problem are executed simultaneously, which may require multiple sequential runs if $M$ is too small to encapsulate them all at once (Appendix E.8.1). The execution time of the subcircuits is based on their compilation onto a linear-chain topology of qubits using a worst-case-scenario swap network [35] with hardware-native gates ($t_{\texttt{Rx}}=40$ ns and $t_{\texttt{ISWAP}}=122$ ns), measurement, and reset times ($t_{\texttt{m+r}}=6~{}\mu$s), all multiplied by $n=5\times 10^{4}$ (shot count), see Appendices E.8.1 and E.4.
Figure 2: The optimal run-time for various classical solvers to match the performance of the QRR* solver with $p=1$, normalized by the run-time of the QRR* (Appendix E.8.1), as a function of the number of variables. The Gurobi data points for the two largest sizes are lower bounds where no solutions matching that of the QRR* have been found in $600$ seconds for most of the problem instances. The quantum run-time component of QRR* is based on $t_{\texttt{Rx}}=40$ ns, $t_{\texttt{ISWAP}}=122$ ns, $t_{\texttt{m+r}}=6~{}\mu$s, and $n=5\times 10^{4}$ shots (Appendix E.8.1). (a) Utilizing a quantum computer with $10^{2}$ qubits. (b) Utilizing a quantum computer with $10^{3}$ qubits. Each data point has been averaged over $1,000$ randomly generated problem instances for each number of variables. The error bars indicate one standard deviation.
@@
We consider two scenarios in which a quantum computer has $M=10^{2}$ (Fig. 2a) and $10^{3}$ (Fig. 2b) qubits, enabling the parallel execution of more QAOA subcircuits. This range corresponds to what is already available and should become available within the next few years, respectively. We show the optimal time $t^{*}_{\textrm{opt}}$ for classical solvers to match the experimental performance of the QRR* with $p=1$, normalized by the run-time of the QRR* itself. In the case of Gurobi, this quantum speedup factor exceeds one for the largest problem sizes, meaning that the QRR* is faster. In fact, for the two largest sizes, we only provide a lower bound corresponding to a $10$-min run-time limit for Gurobi (grayed-out data symbols in Fig. 2), which has failed to match the performance of the QRR* for most of the $1,000$ problem instances (see App. E.9). Regardless, Gurobi provides additional valuable information, such as a bound on how close a solution is to optimality, and it can guarantee the optimality of a solution. The SA and BM solvers are faster than the QRR* at a fixed number of QAOA layers, with the gap converging to a finite value as the problem sizes increase. For SA, we have assumed knowledge of the optimal number of sweeps, which would otherwise need to be searched. Consequently, the quantum speedup factor for SA should be considered a lower bound.
@@
The first path for improving the competitiveness of the quantum solver is to reduce its run-time. The classical steps could be optimized using standard software-engineering techniques and/or algorithmic improvements, and one could imagine the control system of the device estimating expectation values on the fly rather than collecting bit strings that require postprocessing. The execution time on the quantum computer could also be reduced through a more efficient compilation of the quantum circuit onto the hardware-native topology of the qubits, shorter logical gates, and the development of new gates, reducing the compilation overhead. The second path for improving competitiveness is to increase the performance of the QRR*. While additional classical processing is an option, increasing the number of layers $p$ in quantum algorithms remains the cornerstone of more performant quantum algorithms.